OpenStudy (anonymous):

(x+2)^2-(x-3)^2=3x-11, do usual order of operation apply???

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

exponent 1st then add.

OpenStudy (mimi_x3):

x^2 + 4x + 4 -(x^2 - 6x+9) = 3x -11 x^2 + 4x + 4 - x^2 + 6x - 9 = 3x -11 10x - 5 = 3x-11 2x = -6 x = -3

OpenStudy (anonymous):

actually mimi acording to my textbook you are wrong

OpenStudy (mimi_x3):

woops sorry I did it too fast , i will do it again xD

OpenStudy (mimi_x3):

x^2 + 4x + 4 - x^2 + 6x - 9 = 3x -11 10x - 5 = 3x -11 7x = -6 x = -6/7 Is it right :?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

yay ! i just made a stupid mistake before sorry xD

OpenStudy (anonymous):

could we go through you answer together?

OpenStudy (anonymous):

where did you get the 4x from?

OpenStudy (mimi_x3):

Um , okay xD Well, I expanded (x+2)^2 , you know how to expand right ?

OpenStudy (anonymous):

yes (x^2+4)

OpenStudy (mimi_x3):

No, that's not expansion (x+2)(x+2) <-- this is expansion

OpenStudy (anonymous):

oh, i got this

OpenStudy (anonymous):

thanks mimi

OpenStudy (mimi_x3):

yw (: If you still need help, i can help you (:

OpenStudy (anonymous):

lemme try and solve it

OpenStudy (mimi_x3):

Okay (:

OpenStudy (anonymous):

when expanding the 2nd term is is -(x-3)-(x-3) or (x-3)(x-3)???

OpenStudy (mimi_x3):

No, don't do that it will get you confused. Well, when I do it i expand it first so (x-3)^2 = x^2 - 6x + 9 Then , -(x^2 - 6x + 9) -> -x^2 + 6x - 9 Do you get it ?

OpenStudy (anonymous):

i know how to expand.. here is the right way.

OpenStudy (mimi_x3):

LOL, it looks more complicated =/

OpenStudy (anonymous):

hahahaha!

OpenStudy (anonymous):

Loooooooool

OpenStudy (mimi_x3):

Btw, I don't get your way of expanding it xD

OpenStudy (anonymous):

its not simple, you need a great mind to do that and a lot of courage too. :)))

OpenStudy (anonymous):

@mimi, dont get it, I know how to expand, I am wondering if the - sign makes thebefore the second term (x-3) negative

OpenStudy (mimi_x3):

LOL, courage!? you can use an easier way that is called smart xD Put the negative sign outside of the bracket -(x^2 + 6x-9) -> -x^2 - 6x + 9 , Do you get it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

distributing the - sign

OpenStudy (mimi_x3):

yup :D